OpenStudy (anonymous):

A pyramid has a square base with sides of length s. The height of the pyramid is equal to 1/2 of the length of a side on the base. Which formula represents the volume of the pyramid?

OpenStudy (anonymous):

A) \[\frac{ 1 }{ 12 }s ^{3}\] B) \[\frac{ 1 }{ 6 }s ^{3}\] C) \[\frac{ 1 }{ 3 }s\] D) \[3s ^{3}\] E) \[6s ^{3}\]

OpenStudy (anonymous):

@DoomKnight9867 , can you please help me!!(:

OpenStudy (anonymous):

Hmm . . . what grade are you in?

OpenStudy (anonymous):

Sorry I don't know.

OpenStudy (anonymous):

11th, but in taking geometry on plato to make up my credit. my geometry teacher never taought me anything so I failed!

OpenStudy (anonymous):

I'm only 6th grade.

OpenStudy (anonymous):

well, never mind then. thanks for trying!

OpenStudy (anonymous):

@ganeshie8 , can you please help me? (:

OpenStudy (anonymous):

\[Volume = s ^{2}\frac{ h }{ 3 }\]

ganeshie8 (ganeshie8):

Volume of pyramid = \(\large \dfrac{1}{3}Bh\) \(\large B\) = Base area = \(\large s^2\) \(\large h\) = height of pyramid = \(\large \dfrac{1}{2}s\)

ganeshie8 (ganeshie8):

plug the values ^

ganeshie8 (ganeshie8):

Volume of pyramid = \(\large \dfrac{1}{3}Bh = \dfrac{1}{3}(s^2)(\dfrac{1}{2}s) = \dfrac{1}{6}s^3\)

OpenStudy (anonymous):

LOL, i don't know why they go away after asking the question. this problem is with everyone

ganeshie8 (ganeshie8):

@sophiadimeglio

OpenStudy (anonymous):

Thank you so much! sorry, my computer restarted so I had to log back on!

OpenStudy (anonymous):

no problem, i understand your problem. good luck!